HUBBLE SPACE TELESCOPE DAILY REPORT #2263 PERIOD COVERED: 0000Z (UTC) 12/03/98 - 0000Z (UTC) 12/04/98 Daily Status Report as of 338/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Ten Sets of NIC/1/2/3 7963 (Darks: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the detector dark current during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. The observations were completed as planned, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.4 Completed STIS/CCD 7354 (Kinematics of Gaseous Disks in the Nuclei of Liners with Compact Flat Spectrum Radio Sources) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclei of the liner in the compact flat spectrum source galaxy NGC-2911. Many active galactic nuclei channel energetic particles into jets that feed powerful extended lobes of radio emission; others remain quite compact. Both classes of active galaxies are suspected to be powered gravitationally by the accretion of surrounding material onto a massive black hole. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 7622 (WF/PC-2 Cycle 7 Internal Monitor) The WF/PC-2 was used to make a set of Cycle 7 internal monitoring calibration measurements. A variety of internal exposures are obtained in order to provide a monitor of the integrity of the CCD camera electronics in both bays {gain 7 and gain 15}, a test for quantum efficiency in the CCDs, and a monitor for possible buildup of contaminants on the CCD windows. A bias calibration and an internal flat calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2 6967 (Blue Continuum Snapshots of 3CR Radio - Continued Galaxies) The WF/PC-2 was used to make observations of the radio galaxy 3C93.0. Radio galaxies are an important class of extragalactic objects: they represent one of the most energetic astrophysical phenomena; they may be used as probes of their environments; and they are unique probes of the early Universe. The observations were completed as planned, and no problems were reported. 1.7 Completed WF/PC-2 6354 (Measuring Luminosity Evolution in Z = 0.3 Field Galaxies) The WF/PC-2 was used to make snapshot observations of the galactic cluster Gal-Cluster-0597 located at a distance of Z = 0.3. The targets were drawn from a much larger pool of galaxies for which the internal kinematics have been measured from an extensive ground based program. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 7464 (The Evolution of Massive Stars as a Function of Metallicity: Closing the Loop Observationally) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the giant OB class star M-33-OB127 which is located in the M-33 galaxy. The galaxies of the Local Group provide the laboratories for testing how massive stars evolve as a function of metallicity. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Five Sets of NIC/1/2/3 7961 (Flats: Warming Up) The Near Infrared Camera (Cameras One, Two and Three) was used to monitor the flat fields during the interval between the end of science observations and the exhaustion of cryogen and subsequent warming of the dewar to > 100K. These flats will provide a monitor for particulate comtamination and detector lateral position {from the coronagraphic spot and FDA vignetting}. The proposal completed with no reported problems. 1.10 Completed WF/PC-2 6564 (Detailed Morphology of a Lyman-Alpha Galaxy at Z=2.3) The Wide Field Planetary Camera-2 was used to make observations of the Lyman-Alpha emitting galaxy PHL957, which is at a distance of Z = 2.3, in order to study its detailed morphology. The identification and detailed study of galaxies at high redshift remains the most direct path to understanding galaxy formation and evolution. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed Four Sets of STIS/CCD 8062 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations. The observations were completed as planned, and no problems were reported. 1.12 Completed Four Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/MA1/MA2 7657 (STIS Spectroscopic and Imaging Sensitivity, MAMA) The Space Telescope Imaging Spectrograph (MA1 and MA2) was used to make observations of GD71 in order to study the spectroscopic and imaging sensitivity. Sensitivity measurements are done for all primary tilts of the gratings and the prism used with the MAMAs, except for the G140L and G230L gratings. Imaging sensitivity measurements are done with all apertures, including a neutral density filter. A wavelength and a dark calibration image were also taken. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 10 Successful : 10 2.2 FHST Updates: Scheduled : 16 Successful : 16 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. Two TTRs were filed for a total of twenty-one IP data hits on days 337 and 338. A real-time FHST map was commanded at 337/1511Z. The NICMOS error log was dumped at 337/2022Z per ROP NS-11. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH